### v1

# Video Deepfake Abuse: How Company Choices Predictably Shape Misuse Patterns

Abstract.
In 2022, AI image generators crossed a key threshold, enabling much more efficient and dynamic production of photorealistic deepfake images than before. This enabled opportunities for creative and positive uses of these models. However, it also enabled unprecedented opportunities for the low-effort creation of AI-generated non-consensual intimate imagery (AIG-NCII), including AI-generated child sexual abuse material (AIG-CSAM). Empirically, these harms were principally enabled by a small number of models that were trained on web data with pornographic content, released with open weights, and insufficiently safeguarded. In this paper, we observe ways in which the same patterns are emerging with video generation models in 2025. Specifically, we analyze how a small number of open-weight AI video generation models have become the dominant tools for videorealistic AIG-NCII video generation. We then analyze the literature on model safeguards and conclude that (1) developers who openly release the weights of capable video generation models without appropriate data curation and/or post-training safeguards foreseeably contribute to mitigatable downstream harm, and (2) model distribution platforms that do not proactively moderate individual misuse or models designed for AIG-NCII foreseeably amplify this harm. While there are no perfect defenses against AIG-NCII and AIG-CSAM from open-weight AI models, we argue that risk management by model developers and distributors, informed by emerging safeguard techniques, will substantially affect the future ease of creating AIG-NCII and AIG-CSAM with generative AI video tools.
copyright: none
Content notice: this paper discusses AI-generated non-consensual intimate imagery and child sexual abuse material.
1. Introduction
In 2022, AI image generation crossed a critical threshold, with models like DALL-E 2 (ramesh2021) and Stable Diffusion 1.x (Rombach2021) making the creation of photorealistic synthetic images much more efficient, dynamic, and accessible than ever before. In particular, Stable Diffusion’s 1.0’s open release in August 2022 made access to these capabilities widely available, allowing large numbers of users to create photorealistic images with minimal technical expertise, very little time, and no specialized data (petsiuk2022human). This enabled positive uses of image diffusion systems for recreation, graphics, and art. However, the accessibility of these new systems enabled unprecedented levels of misuse for AI-generated non-consensual intimate imagery (AIG-NCII), including AI-generated child sexual abuse material (AIG-CSAM). Since 2022, AIG-NCII and AIG-CSAM images have surged (securityhero2023; iwf2023ai-csam; iwf2024; NCMEC2024). For example, a report from ActiveFence estimated that the number of “threads related to the creation of [AIG-]NCII depicting private individuals rose…by 400%” between 2022 and 2023 (activefence2023). New image diffusion models (principally Stable Diffusion and Flux models, schneider2024image; hawkins2025deepfakes) and the infrastructure built around them (e.g., Chai2022) demonstrated how a small number of models capable of producing realistic NSFW deepfakes can drive misuse patterns (see Section 2).
Figure 1. The supply chain for open-weight AI models capable of creating non-consensual intimate video deepfakes. Models flow from developers through distribution platforms to modifiers, who create specialized variants that power user-facing applications (e.g., undressing applications). Individual actors with technical expertise can also directly download models from distribution platforms and create AIG-NCII locally. The modification and redistribution cycle (highlighted in orange) shows how models with openly available weights can undergo multiple rounds of modification and be re-uploaded to model distribution platforms. Developers and distribution platforms serve as critical bottlenecks. Scale indicators show the rough number of actors at each stage.
The image generation transformation of 2022 offers precedent for understanding current developments in video generation, where similar capability thresholds are now being crossed. Currently, in 2025, video generation models are undergoing an analogous revolution. Systems such as OpenAI’s Sora (openai2024), Google’s Veo (google2025), and Runway’s Gen-4 (runway2025) can be used to produce convincing videorealistic content. Meanwhile, as we will discuss in Section 3, a small number of open-weight models are emerging as the dominant tools for NSFW video generation, including Wan2.x. Variants of these models specialized for NSFW content are widely shared across several key online distribution platforms, including CivitAI, making these platforms critical gatekeepers for access to these capabilities.
This paper examines how developer and distribution platform [^1] This paper uses “distribution platform” to refer to model and dataset distribution platforms like CivitAI, Hugging Face, and GitHub. These are distinct from further downstream platforms like social media sites or adult sites where AIG-NCII is distributed. choices can shape patterns of video deepfake misuse. By analyzing the history of image generators, current AI videos shared online, discussions on online forums, and academic literature on model safeguards, we present three key findings:
  1. (1)
    NSFW AI video content online disproportionately stems from a small number of models, including Wan 2.x, whose variants are largely distributed through a small number of online distribution platforms, including CivitAI (Section 3).
  2. (2)
    Various technical strategies can impose significant barriers to using these models for AIG-NCII and AIG-CSAM (Section 4).
  3. (3)
    Friction-based mitigations meaningfully reduce harm even when perfect prevention is not possible (Section 5.1).
Based on this evidence, we draw two conclusions about how company choices shape downstream misuse patterns:
  1. (1)
    AI developers who (a) train videorealistic video generation models without effective filtering of NSFW content, (b) omit sufficient post-training safeguards, and (c) release them with open weights foreseeably contribute to mitigatable downstream harm.
  2. (2)
    Model distribution platforms that do not proactively monitor models for misuse and take down ones designed for AIG-NCII and AIG-CSAM amplify these harms.
We conclude that the decisions made by AI video generator developers and distributors in the coming months will have the potential to influence misuse patterns into the future, making current attention to risk mitigation important for reducing future harm.
2. The History of Photorealistic NSFW Image Generation and the 2022 Transformation
2.1. Methods Prior to 2022
Prior to the rise of generative AI, convincing synthetic imagery depicting real humans was well-precedented. Early methods relied on manual image editing tools such as Photoshop, which could enable skilled users to stitch together images of different humans’ faces and bodies (eggestein2014fighting). In the 2010s, progress in research and development of generative AI added image and video processing models to the toolkit (monaghan2017impact; gieseke2020new; henry2020image). From 2017 onward, there was a notable rise in the availability of both “AI nudification” tools (for both images and video) and cultural awareness of them (kobriger2021out; wagner2019word). The gradual rise of tools that could enable AIG-NCII was driven, in part, by the development and release of open software tools from internet communities dedicated to synthetic NSFW content generation (monaghan2017impact; winter2020deepfakes).
Prior to 2022, despite it being possible to generate photorealistic NSFW images and videos of real individuals, tools were limited by skill requirements, a lack of dynamism, inefficiency, and/or inconsistent realism. For example, Photoshop required significant time and experience. Meanwhile, early generative adversarial networks (GANs) suffered from instability, mode collapse, and limited coherence, producing images that were often easily distinguishable from authentic photographs (saad2024). Prior to 2022, it was particularly challenging for GANs to consistently and dynamically generate realistic images. Models like AttnGAN and DM-GAN performed poorly on standard benchmarks and struggled with complex compositional understanding (xu2017; zhu2019). The most accessible approaches, such as VQGAN+CLIP combinations, required extensive technical expertise, complex prompt engineering, and produced imagery with characteristic visual artifacts that marked them as clearly synthetic (crowson2022; steinbruck2022). Even OpenAI’s DALL-E 1 diffusion model, while demonstrating impressive conceptual capabilities, operated at low resolution and was not publicly accessible (ramesh2021).
2.2. The DALL-E 2 Release
DALL-E 2 ushered in the modern age of dynamic photorealistic image generation. Released by OpenAI in April 2022, DALL-E 2 marked a leap forward in the ability of AI tools to consistently and dynamically produce approximately photorealistic results at high resolution (1024×1024) from simple text prompts (ramesh2022dalle2). OpenAI kept the model behind a closed API and reported implementing technical safeguards to prevent misuse, including filtering NSFW training data; internal testing; independent red team evaluations; prompt filtering to block explicit content and celebrity names; and filtering generated images for NSFW content. Furthermore, OpenAI required user registration with email verification and stated intentions; implemented user suspension systems for suspicious activity; and performed continuous backend updates to address emerging issues (dalle2modelcard). Subsequently, some researchers have found that these safeguards were imperfect, allowing for the production of some NSFW content (yang2024sneakyprompt). However, to the best of our knowledge, there is no evidence of successful, scalable uses of DALL-E 2 for the creation of individualized AIG-NCII.
2.3. Stable Diffusion’s Release and Downstream Misuse
Stability AI released Stable Diffusion 1.0 without effective mitigations against AIG-NCII misuse. Stable Diffusion 1.0 was released in August 2022 with open weights. It was released with a license, a user agreement, beta testing, and an NSFW classifier to block NSFW content. However, these measures proved to offer no substantial barrier to misuse. Usage terms are effectively unenforceable for an open-weight model; the beta testing did not reportedly involve red-teaming for AIG-NCII risks (stable_diffusion_launch_2022); and the NSFW content classifier could be trivially disabled by users who downloaded the system. [^2] The Stable Diffusion safety filter was also vulnerable to attacks (Rando2022). Meanwhile, Stable Diffusion 1.x models were trained on weakly-curated internet data (the LAION-5B dataset), which contained a substantial amount of NSFW content and CSAM, (birhane2311into; thiel2023). Meanwhile, Stability AI did not publicly report on any red teaming for NSFW capabilities or AIG-NCII risks. As a result of these ineffective safeguards, derivatives of Stable Diffusion 1.x models empirically became effective tools to generate AIG-NCII (schneider2024image; hawkins2025deepfakes).
Community-developed tools greatly increased access to AIG-NCII creation in subsequent months. Despite the photorealistic image generation power of models like Stable Diffusion 1.x, they were not particularly dynamic or accessible tools by themselves. For example, initial misuse of them for AIG-NCII was largely focused on generating content of celebrities. The final key breakthrough toward widespread (mis)use was the introduction of software and applications that greatly increased dynamism and accessibility (wagner2025perpetuating). For example, DreamBooth, released in September 2022, enabled personalization of diffusion models using as few as three to five reference images (Chai2022). Using DreamBooth, users could “teach” models to associate unique identifiers with specific individuals, then generate novel synthetic content featuring those individuals in any described scenario. Meanwhile, other techniques, such as inpainting, outpainting, and stitching, [^3] Inpainting refers to having a diffusion model construct a customized synthetic version of an image region. Outpainting refers to having a diffusion model synthetically expand an image beyond its original border. Stitching refers to having a diffusion model realistically blend two separate images together (such as inserting one person’s head onto another person’s body). allowed users to produce NSFW content of specific subjects from SFW images (shani2021). These innovations meant that a person’s social media photos, school yearbook pictures, or family photos were sufficient for creating personalized photorealistic AIG-NCII images, expanding potential targeting from public figures to virtually anyone with minimal digital presence. Furthermore, community development quickly produced models further fine-tuned on NSFW data, user-friendly interfaces, pre-configured workflows, and optimized implementations accessible to non-technical users. This decentralized supply chain of these technologies creates a “malicious technical ecosystem” (ding2024malicious) that has enabled dozens of “AI nudifier” applications, which allow nontechnical users to create AIG-NCII within minutes (gibson2025analyzing; alexios_mantzarlis_ai_2025).
2.4. Misuse Patterns Following Technical Breakthroughs
A surge in personalized AIG-NCII followed the NSFW image generation breakthroughs of 2022. Within months of Stable Diffusion’s release, dedicated communities formed around generating non-consensual content, sharing techniques for circumventing safety measures, and distributing specialized model configurations optimized for explicit content generation. The vast majority of which depicts women and girls (securityhero2023; laffier2023deepfakes; my_image_my_choice_deepfake_2024; wagner2025perpetuating). Reports involving generative AI to the National Center for Missing and Exploited Children’s CyberTipline surged from 4,700 in 2023 to 67,000 in 2024, a 1,325% increase (NCMEC2024). [^4] These reports include AI-generated CSAM as well as other forms of child sexual exploitation involving generative AI, such as grooming guides and nudify applications. Analysis by Graphika found that 34 synthetic AIG-NCII providers received over 24 million unique visitors to their websites in September 2023, with referral link spam for these services increasing by more than 2,000% on platforms since the beginning of 2023. Graphika also found 52 Telegram groups used to access AIG-NCII services containing at least 1 million users as of September 2024 (graphika2023). Meanwhile, Activefence reported that “threads related to the creation of AIG-NCII depicting private individuals rose in the same period by 400%” (activefence2023). Ultimately, these developments have led to unprecedented patterns of AIG-NCII misuse. For example, in early 2025, law enforcement identified and made dozens of arrests connected to an international online community with hundreds of members dedicated to AIG-CSAM (europol2025operation).
3. The AI Video Generation Ecosystem in 2025
“This is going to really have some wild implications for people who share their photos. Throw a pic up on Instagram, and in ten minutes people can make fairly realistic video clips of you performing sex acts.”
- A Reddit user commenting under a series of NSFW videos created from an SFW image (link omitted due to NSFW content).
Next, we analyze current patterns in the use of AI video generators for creating NSFW content, observing clear parallels with image models in 2022 and 2023. However, there is limited open research on NSFW video content generation because of ethical barriers, legal barriers, and social taboos. Given the rapid pace of image-generation technology development and the “underground” nature of NSFW generation that may not be captured by formal literature, we also analyze online platforms, discussion forums, publicly accessible applications, and user communities to understand usage patterns.
3.1. Technical Capabilities and Accessibility
Video generation models in 2025 are crossing an analogous capability threshold to what enabled widespread AIG-NCII in 2022. Current systems, including Google’s Veo, OpenAI’s Sora, and Runway’s Gen-3, can often produce temporally coherent synthetic videos with quality approaching broadcast standards (openai2024; google2025; runway2025). Motion fidelity and facial consistency have reached levels where synthetic content sometimes becomes difficult to distinguish from authentic footage. [^5] Subjectively, we attest that a small fraction of the NSFW video deepfakes that we have seen online have appeared to us to be entirely indistinguishable from genuine video.
Model distribution platforms, principally CivitAI, facilitate the spread and usage of AI video models for NSFW content. Model distribution platforms provide infrastructure to help users download models and access detailed workflow documentation (gorwa2024moderating). By a large margin, the most prominent distribution platform for distributing models that specialize in NSFW content is CivitAI (wagner2025perpetuating; hawkins2025deepfakes; wei2024exploring; civitai2024transparency), who reported 446K LoRA adapters trained and 600M images/videos downloaded using their site in 2024. Community-developed workflows (e.g., from https://civitai.com/articles) further increase the accessibility of NSFW model generation capabilities by providing instructions, workflows, optimized settings, etc. that enable users to generate custom NSFW videos in minutes (fusionxcivitai2025).
3.2. Current Patterns and Methods for NSFW Video Generation
Figure 2. Which open-weight video generation models are the most disproportionately used to create NSFW content online? We analyze model search hits on subreddits (left), model search hits on CivitAI and a CivitAI model archive site (middle), and video search hits on CivitAI (right). In each analysis, we report the SFW market share, the NSFW market share, and the NSFW/SFW market share ratio. The first two columns of each grid sum to 100%. Some models, including Wan2.x, stable-video-diffusion, HunyuanVideo, and LTX-Video are disproportionately used to generate NSFW content.
A small number of open-weight models account for the majority of model variants and videos across prominent online NSFW AI video communities To understand current usage patterns, we analyzed model search hits and tags across web communities. For this analysis, we assume that a model’s usage for generating NSFW content in general is an informative proxy for its usage in creating AIG-NCII (see Appendix A for further discussion). First, we identified 10 popular open-weight video generator model families online. [^6] We consider a model family to refer to a named and versioned set of models from a single developer, including all releases and derivatives. For example, we considered Wan-AI/Wan2.1-T2V-14B, Wan-AI/Wan2.2-TI2V-5B, and Comfy-Org/Wan_2.2_ComfyUI_Repackaged to all be from the same Wan2.x model family. We did this by exhaustively searching for all general-purpose text-to-video, image-to-video, and video-to-video models on Hugging Face from 2024 or later that had more than 10,000 downloads in the past month (as of October 8, 2025). Next, we identified a set of web platforms that allowed us to search and count AI video models specialized for SFW and NSFW content and AI videos containing SFW and NSFW content. Full details of our site selection process are in Appendix A. We ultimately selected six subreddits, CivitAI (the world’s most popular community platform for distributing AI models fine-tuned for generating NSFW content (civitai2024transparency)), and a website dedicated specifically to archiving models and content from CivitAI to evade moderation.
To quantify content distribution across platforms, we used a systematic search procedure. For subreddits, we searched each model name across three SFW (r/aiArt, r/StableDiffusion, r/aivideo) and three NSFW (names omitted) subreddits, recording the number of search results returned. For SFW subreddits, we performed these searches using NSFW filters. For CivitAI and the CivitAI model archive website (name omitted), we conducted searches for each model name under two conditions: with the platform’s NSFW filter disabled (capturing total content) and enabled (capturing only SFW content), deriving NSFW counts by subtraction. For video-specific analysis on CivitAI, we used a dual approach. First, we identified NSFW videos by searching with model filters applied and three explicit search terms (“nude,” “NSFW,” “naked”); second, we captured SFW video content by searching with model filters and an empty search query while enabling the NSFW filter. For each search, we recorded the total number of results reported by the platform’s search functionality. From these raw counts, we calculated three metrics for each model-platform combination: SFW market share, NSFW market share, and an NSFW/SFW market share ratio, where values exceeding 1.0 indicate disproportionate usage for NSFW content (Figure 2).
As shown in Figure 2, different models’ SFW and NSFW market share differ across platforms. We order models based on their average NSFW/SFW market share ratio across all three analyses. Some models have disproportionately high market share for NSFW content. In particular, this includes Wan2.x, stable-video-diffusion, HunyuanVideo, and LTX-Video.  [^7] For comparison, we also analyzed xAI’s Grok Imagine model, which is fairly unique among frontier closed-weight AI video generators because its deployers have intentionally enabled its API to produce NSFW video content (maiberg2025grok; rainn2025grok). We analyzed the prevalence of videos from Grok Imagine on Reddit using the same methodology as with the open models. Despite accounting for 12.7% of NSFW discussions in our Reddit sample, Grok’s NSFW-to-SFW ratio was only 0.58, substantially lower than open-weight models like Wan2.x (1.39) and Stable Video Diffusion (1.04).
4. Mitigating Harms from Open-Weight Models: the Role of Developers and Distributors
4.1. How can Developers Reduce Risks?
Existing literature on AI safeguards and risk management establishes that several techniques can significantly mitigate harmful uses of open-weight AI models (casper2025open). Here, in line with casper2025open, we focus on risk mitigation techniques for open-weight models that cannot be easily circumvented by users. [^8] In addition to risk mitigation techniques that we present in Section 4.1, data provenance methods also serve an important role. While these techniques do not directly mitigate risks, they facilitate a broader understanding of the ecosystem. See bengio2025international and casper2025open. For example, while external safety filters can be valuable to deploy alongside open-weight models (e.g. agarwal2025cosmos), they can be trivially disabled and are empirically weak safeguards. Crucially, comprehensive AIG-NCII risk mitigation strategies for open models must account for how fine-tuned derivatives of open-weight models, rather than the original models themselves, are empirically a major contributor to misuse (simonovich2024nytheon; casper2025open). For example, CivitAI hosts thousands of fine-tuned low-rank adapters designed to make video generation models more consistent, dynamic, or specialized in their ability to produce NSFW content.
4.1.1. Training Data Curation
Current evidence suggests that data filtering is a useful intervention for mitigating harmful content generation capabilities. Filtering content from training data related to harmful topics helps developers train models with minimal knowledge of such content (maini2025safety; o2025deep; nichol2021glide). While the relationship between training data filtering, model capabilities, and resistance to harmful fine-tuning is not yet fully understood in diffusion models, filtering NSFW content from training data is recognized as a key defense against AIG-NCII (NIST_ID012_2024; NIST_ID012_2024). A model trained on NSFW-filtered data can always be taught to generate NSFW content after enough fine-tuning, but removing NSFW content from training data substantially increases barriers to misuse by requiring users to fine-tune new NSFW capabilities into a model before it can be misused (o2025deep; nichol2021glide). Dataset curation can also be an important point of harm remediation in machine learning research, as some datasets used in open science contain non-consensually obtained datasets of NSFW content (Cintaqia2025Stop).
Some approaches to filtering NSFW data can be insufficient to prevent misuse, as suggested by Wan2.x. While filtering reduces the ease of misuse, it can be insufficient. For example, in the Wan 2.1 technical report, Alibaba reported to “systematically evaluate and filter inappropriate content based on computed NSFW scores in all training data” with no additional details provided. However, given the model’s documented popularity for NSFW content generation (see Section 3.2), this filtering appears not to have been a strong safeguard. Due to a lack of additional public information about how Alibaba filtered NSFW content, it is not possible to critically assess the effectiveness of their specific approach to NSFW data filtering. [^9] We contacted Alibaba, asking for comment and additional details, but received no reply. However, we can turn to image diffusion models for a notable empirical success story of data filtering.
Stable Diffusion 2.0 provides empirical evidence for the effectiveness of filtering NSFW data as a safeguard. Stability AI released Stable Diffusion 2.0 in November 2022, reporting that “Stable Diffusion 2.0 delivers several big improvements and features versus the original V1 release” (StableDiffV2Release2022). Stable Diffusion 2.0 was trained on “an aesthetic subset of the LAION-5B dataset…further filtered to remove adult content using LAION’s NSFW filter” (StableDiffV2Release2022). Subsequently, the NSFW content generation community has empirically found Stable Diffusion 2.0 to be unhelpful for generating NSFW content. For example, we identified various (SFW) Reddit posts reporting “We all know by now that SD 2.0 is pretty poor for celebrities, artists, and NSFW content,” that Stable Diffusion 2.0 was “abandoned because it can not generate any NSFW content,” and “Well if you’re an overly sensitive prude who has anxiety attacks if you see some [redacted], SD 2.0 is great because it’s so heavily censored.” We confirmed this using analysis of search hits on CivitAI (Figure 3). Across three search terms (“nsfw”, “nude”, and “naked”), Stable Diffusion 1.x models and derivatives had dramatically more NSFW content than Stable Diffusion 2.x models: 179,345+ total hits compared to only 128, resulting in a ratio of 1,401×. [^10] The “nsfw” search for Stable Diffusion 1.x exceeded CivitAI’s display limit of 100,000 results, so the actual difference is likely substantially larger. Meanwhile, the ratio of NSFW to NSFW fine-tuned models is over twice as large for Stable Diffusion 1.x.
Figure 3. Stable Diffusion 2.x models offer an empirical example of models trained without NSFW data being used less for generating NSFW content. On CivitAI, Stable Diffusion 1.x has over 1,000x more tagged NSFW images and over 2x more NSFW fine-tuned models compared to Stable Diffusion 2.x. (Left) NSFW image searches on CivitAI show Stable Diffusion 1.x dominates with 100,000+ results for ”NSFW” and tens of thousands for other terms, while Stable Diffusion 2.x returns fewer than 100 results per term. (Right) Model count analysis reveals Stable Diffusion 1.x has 37,075 NSFW models compared to Stable Diffusion 2.x’s 186 models (199x difference), with NSFW/SFW ratios of 0.49 and 0.20, respectively. Model counts obtained by searching CivitAI with base model filters for each Stable Diffusion version. Stable Diffusion 2.0’s substantial reduction in both NSFW content and models offers evidence that training data filtering can effectively reduce a model’s utility for NSFW generation without eliminating broader capabilities.
4.1.2. Fine-Tuning, Unlearning, and Anti-Tampering
Post-training safeguards can significantly increase barriers to misuse. After a diffusion model is trained, post-training safeguards can offer a second line of defense against misuse of NSFW content generation capabilities. In particular, developers can utilize “machine unlearning” methods to suppress NSFW-generation capabilities. In recent years, researchers have introduced and tested a number of approaches to unlearning specific capabilities from diffusion models, including NSFW content (zhang2024defensive; fuchi2024erasing; zhu2024choose; wu2025erasing; zhang2025concept; wu2025unlearning; lu2025concepts). To date, the bulk of research on these methods has focused on image diffusion models, but the same algorithms are equally applicable to video diffusion models (e.g., liu2024unlearning). While useful, most current unlearning methods are brittle and can be partially circumvented by adversarial prompts or adversarial few-shot fine-tuning (sharma2024unlearning; suriyakumar2024unstable; zhang2024generate; george2025illusion; yu2025forgetme). However, there is also emerging research on diffusion model unlearning methods that resist relearning (gao2024meta; li2025towards; abdalla2025gift). For example, abdalla2025gift introduced an unlearning method that caused diffusion models to resist up to 2,500 steps of fine-tuning on NSFW content while also preserving the model’s ability to learn from benign images. Overall, research is still ongoing to develop more practical and relearning-resistant unlearning techniques for video diffusion models. Nonetheless, existing post-training methods for suppressing NSFW generation capabilities offer a barrier to misuse without significantly affecting the SFW capabilities of a diffusion model.
4.1.3. Evaluations
Evaluating open-weight models for misuse potential, particularly under fine-tuning threats, is needed to meaningfully assess real-world misuse risks. Evaluations (both internal and external) of frontier AI systems’ capabilities are increasingly recognized as a key pillar of AI risk management frameworks. Safety evaluations of frontier systems are most commonly conducted using input-output access in which evaluators give models queries and analyze their responses. For closed weight models, input-output evaluations emulate the real-world misuse threats that models will be exposed to. However, for open-weight models, input-output evaluations do not fully account for the risks posed by simple downstream modifications (casper2024black; che2025model; casper2025open). While it is intuitive that evaluating the safety of open-weight models requires adversarial fine-tuning, there is limited precedent for open-weight models to be evaluated under harmful fine-tuning attacks. For example, while some of the developers we analyzed in Section 3 report on evaluations for NSFW capabilities (see Table 1), zero mention evaluations under fine-tuning.
4.1.4. Staged Deployment
Staged release strategies enable monitoring and patches prior to an open-weight release. Model deployment is not a binary between fully closed and fully open. Different deployment strategies offer different trade-offs between openness and security. When a company has the resources to implement them, these options allow model developers to gradually increase access to a model before choosing to fully release it with open weights (solaiman2023gradient). For example, a developer could first provide API query access to a model to study its input-output uses. Second, they could offer access to a fine-tuning API to study risks from fine-tuning. Third, they could release only some layers of the model and enable access to the others through a collaborative inference API (ren2023survey). These strategies allow developers to monitor how real users interact with their systems, identify emerging misuse patterns, and refine safety measures before a full open release (casper2025open). For example, observations following a staged deployment could help the developer identify a need for additional mitigations, such as unlearning (Section 4.1.2).
4.2. Developers Rarely Report on Mitigations
Most open-weight model developers do not report on efforts to safeguard their models against AIG-NCII misuse. Above in Section 3.2, we observe that some models, including Wan2.1, stable-video-diffusion, HunyuanVideo, and LTX-Video are disproportionately used for NSFW content. To better understand why this might be the case, we analyze technical reports from developers for mentions of the mitigation strategies discussed above in Section 4.1.111111Our analysis was limited to reports available online in English. However, some of the developers in Table 1 are based in countries where English is not the official language. Further details may be available in non-English languages. We summarize results for the same ten models as in Section 3.2 in Table 1. With the exception of Cosmos (agarwal2025cosmos), developers consistently report very little about AIG-NCII risks or risk mitigations. And even when information is presented, it is often with very little detail.
Training Data
Curation
Fine-Tuning,
Unlearning, &
Anti-Tampering
Evaluations
Staged
Deployment
Mention of NCII
in Model License
Wan-AI/Wan2.x
wan2025wan
<1<1 Paragraph No Mention No Mention No Mention No Mention
Stable Video Diffusion
blattmann2023stable
No Mention No Mention No Mention No Mention $\geq 1$ Paragraph
HunyuanVideo
kong2412hunyuanvideo
No Mention No Mention No Mention No Mention <1<1 Paragraph
LTX-Video
hacohen2024ltx
No Mention No Mention No Mention No Mention No Mention
SeedVR2
wang2025seedvr
No Mention No Mention No Mention No Mention No Mention
CogVideoX
yang2024cogvideox
<1<1 Paragraph No Mention <1<1 Paragraph No Mention No Mention
AnimateDiff-Lightning
lin2024animatediff
No Mention No Mention No Mention No Mention $\geq 1$ Paragraph
Stable Virtual Camera
zhou2025stable
No Mention No Mention No Mention No Mention $\geq 1$ Paragraph
Cosmos
agarwal2025cosmos
<1<1 Paragraph $\geq 1$ Paragraph No Mention No Mention <1<1 Paragraph
Mochi 1
genmo2025mochi1
<1<1 Paragraph No Mention No Mention No Mention No Mention
Table 1. Popular open-weight AI video generator developers rarely report on AIG-NCII risks and mitigations in model technical reports. When risks or mitigations are discussed, it is often with very little detail.
The lack of reporting from developers makes it challenging to empirically study practical AIG-NCII risk mitigations for open-weight video models. While prior research discussed above in Section 4.1 has established the effectiveness of some safeguards for open-weight models, the lack of reporting from developers on whether/how these techniques were used makes it difficult to assess their usefulness and practice. We reached out to all 10 developers for comment and additional information, receiving only one brief reply from the developers of CogVideoX, clarifying that an unspecified quantity of NSFW data filtering and capability evaluation was used. Overall, based on usage patterns observed in Section 3 and Figure 2, there is evidence that some models were ineffectively safeguarded against NSFW capabilities, particularly Wan2.1, stable-video-diffusion, HunyuanVideo, and LTX-Video. This, combined with a consistent lack of reporting may be indicative of an industry pattern of omitting established mitigations for misuses of NSFW AI-generated video content. However, due to insufficient information, it is not possible to understand current industry practices with precision.
4.3. How Can Model Distribution Platforms Reduce Risks?
Model distribution platforms are key to the accessibility of harmful model capabilities. Thus far, we have primarily focused on video model developers because development is a key point of influence and leverage in the AI ecosystem (williams2025regulating). However, while developers are responsible for creating models, distribution platforms are principally responsible for their widespread accessibility (gorwa2024moderating).
Distribution platforms can amplify misuse by hosting and distributing models designed for AIG-NCII. Many early image-generation tools for of generating AIG-NCII were hosted and distributed on GitHub, one of the most popular developer platforms in the world. In 2023, the National Center on Sexual Exploitation reported that GitHub had been hosting open-source repositories that enable AIG-NCII generation. This included DeepFaceLab, which “contains direct links to the most prolific sexual deepfake abuse website in the United States: Mr.DeepFakes,” as well as repositories for DeepNude and Unstable Diffusion (patrick_trueman_esq_github_2023).
In the past several years, generative AI model distribution has increasingly shifted to CivitAI and Hugging Face. In May 2025, following new regulations and pressure from payment processors, CivitAI banned models designed to recreate real individuals’ likenesses, which, according to 404 Media, dealt a “major blow to the nonconsensual AI porn ecosystem” (civitai404media; civitai404media1). However, the deepfake community subsequently organized a coordinated effort to archive over 5,000 of these models and re-upload them to Hugging Face (maiberg2025). Recently, Hugging Face has been found to host models specifically designed for producing AIG-NCII of celebrities, including individuals who were minors when source material was created (hashim2025; maiberg2025). These models violated Hugging Face’s stated content policy prohibiting “sexual content created without explicit consent” and “underage nudity or any sexual content involving minors” (hfcontentpolicy2025). Nonetheless, maiberg2025 found that Hugging Face was slow to take these models down, allowing some to remain accessible for nearly five months despite multiple requests. Even after removal events, some violating models remained accessible (hashim2025). This pattern illustrates how a lack of proactive enforcement of anti-AIG-NCII policies on model distribution platforms can make harmful models significantly more widely accessible.
5. Discussion
5.1. Are AIG-NCII Harms from Open Video Generators Unmitigatable?
In this paper, we have argued that risk management practices related to data curation, safeguards, evaluations, release, and platform moderation (Section 4) will have a significant effect on the production of AIG-NCII. But are mitigations effective if they are not used universally? A possible argument against prioritizing risk mitigation is that the value of one company’s efforts can be undercut by another company that does not prioritize risk mitigation. We term this the “cat-out-of-the-bag” argument. With models such as Wan2.x already being released and distributed, is the cat out of the bag? Does their existence (or the future existence of other models like them) nullify the value of future models and distribution platforms being safeguarded to mitigate AIG-NCII? We argue it does not.
The “cat-out-of-the-bag” argument overlooks harm reduction principles and incorrectly characterizes technology adoption dynamics. The ‘cat-out-of-the-bag’ reasoning is not in line with established approaches to risk management and misidentifies how technology proliferates. Mitigations matter. For example, digital piracy provides one clear parallel: despite the theoretical availability of pirated content online, coordinated enforcement efforts by search engines, authorities, and companies to reduce the accessibility of pirated content online have had a measurable impact. For example, smith2023piracy concluded that removing links to pirated content from search results increases legal content sales by 11-14%, and (danaher2017) concluded that search engines that demote piracy sites in results shift user behavior toward legal consumption. Even though determined users can still find pirated content with sufficient effort, increasing barriers to accessing it meaningfully reduces piracy by increasing search costs and directing casual users toward legitimate alternatives. Reducing the accessibility of harmful generative AI capabilities may be particularly helpful for reducing AIG-NCII given that a large number of perpetrators are minors themselves (alexander2025deepfake).
Risk mitigation operates through aggregate effects, creating systemic benefits even when circumvention remains possible. Safety measures implemented by major platforms and developers create systemic benefits, including reducing total explicit content volume by avoiding accidental harms, preventing unsophisticated misuse, deterring costly circumventions of safeguards, preventing normalization by keeping harmful applications marginalized rather than mainstream, and enabling legal frameworks that can target deliberate safety circumvention (ec2019).
Technology adoption follows network effects that amplify early choices made by major actors. The first widely available implementations establish user expectations, development priorities, and community norms that influence subsequent development. For example, the LAION dataset was introduced in March 2022 (laion2022laion5b), was found to contain CSAM in December 2023 (thiel2023), and was modified to mitigate CSAM content in August 2024 (laion2024relaion5b). Nonetheless, prior to August 2024, the original LAION dataset was widely used to train generative systems, including Stable Diffusion models (Salvaggio2024), that are still commonly used today. When major developers and platforms implement robust safety measures, they also create incentives for compatible approaches across the ecosystem. Conversely, early releases without safety measures normalize applications and create communities focused on circumvention rather than beneficial use. As such, the “cat-out-of-the-bag” argument offers a pretext to engage in a race to the bottom.
5.2. Policy Influences Company Incentives
This paper has focused on how the choices made by developers and model distribution platforms will affect downstream misuse patterns for AI video generators. However, it is key, especially for policy audiences, to recognize that these choices are not made in a vacuum, but amidst external commercial, regulatory, and liability incentives. Diffusion model generators and distribution platforms have commercial incentives to deploy and share models as quickly as possible, but regulatory requirements, liability, and transparency can have a significant counteractive effect to incentivize risk mitigation (bommasani_considerations_2024; williams2025regulating; smith2024liability). When considering regulation related to managing AI risks– especially from open-weight models, policymakers should recognize that some types of risk management techniques, such as staged deployment, may be more onerous for small developers and academics than for larger companies.
Some laws on AIG-NCII have been written to regulate content and models. Some regulatory approaches have aimed to influence the AIG-NCII ecosystem by targeting downstream non-consensually distributed AIG-NCII, such as the federal TAKE IT DOWN Act in the USA (sen_cruz_ted_r-tx_take_2023). Other approaches can also focus on models themselves. For example, Arkansas Act 827, enacted April 2025, states that “the Attorney General may institute a civil action on behalf of the state against a provider or developer of image generation technology that was used to create deepfake visual material” if “the provider or developer of the image generation technology did not have reasonable safeguards in place to protect against the generation of deepfake visual material” (state_of_arkansas_act_2025).
Transparency requirements invite external scrutiny. Another key factor in the incentives of developers and model distributors is transparency (widder2022limits). Detailed information about risk management measures gives external stakeholders the ability to study their mitigations. A low-transparency state of affairs, as illustrated in Table 1, can be expected by default because selective transparency helps companies avoid unwanted forms of scrutiny (ananny2018seeing). For example, the lack of information about AIG-NCII risk mitigations from developers (see Table 1) impedes work like ours from directly identifying specific links between company choices and downstream harms. Absent external incentives, a low transparency regime will tend to self-reinforce by inviting increased levels of scrutiny and criticism to developers who voluntarily choose to be more transparent.
5.3. Weighing Competing Considerations
Balancing the benefits and risks of open-weight models: As we have discussed, models with openly available parameters are a key factor and key driver behind the creation of AIG-NCII and AIG-CSAM (iwf2023ai-csam; widder2022limits). However, this does not imply open-weight models are net harmful. They are also key for positive uses, including the de-concentration of power and the facilitation of open research on safety (kapoor_societal_2024; bommasani_considerations_2024; francois_different_2025). More pragmatically, powerful open-weight models are increasingly influential (bhandari2025forecasting) and likely inevitable. In recognition of this, we have emphasized a harm mitigation approach for safeguarding open-weight models in Section 4, arguing that substantial reduction of AIG-NCII harms can be achieved without prohibiting powerful open-weight models.
Balancing the benefits and risks of NSFW AI Content: AIG-NCII is a serious form of image-based sexual abuse that results in psychological, financial, and reputational harm, disproportionately borne by women and girls (american_sunlight_project_deepfake_2024; henry_ajder_state_2019; flynn_deepfakes_2022). Prior works have argued that the current NSFW generative AI ecosystem is characterized by non-consensual practices (gibson2025analyzing; ding2024malicious). However, AI models that produce NSFW content could also have applications based on consensual and ethical use. (geeng2025co). Work to build infrastructure for more positive uses of NSFW AI content between adults would require consent-focused data collection, conscientious deployment, and concerted effort to reduce AIG-NCII and AIG-CSAM risks (geeng2025co; Cintaqia2025Stop).
5.4. Limitations
Studying the NSFW content generation ecosystem remains challenging, especially in a low-transparency regime. Our findings are necessarily based on observable reporting, usage patterns, and analysis of discussions in online forums. For example, with little reporting from developers about how they mitigate risks, we have not been able to directly analyze the relationship between mitigations and downstream harms. Future technical work should continue to study the NSFW AI image and video generation ecosystem toward a better understanding of the relationship between upstream decisions and downstream harms. However, absent public pressure or regulatory incentives for companies to meaningfully report about risk assessments and risk mitigation, the empirical study of AIG-NCII and AIG-CSAM harms may continue to be challenging.
Acknowledgments
We are grateful to Markus Anderljung for discussion and feedback.
Appendix A Methodological Details
Search for online platforms: Using web searches, analysis of internet fora, and discussions in past literature, we identified over 50 websites, social media sites, apps, platforms, and Discord servers. We ultimately finalized on a selection of subreddits, CivitAI, and a site dedicated to the archival of CivitAI models because they uniquely allowed us to analyze the number of search hits or tags for SFW and NSFW video models or AI videos. The rest of the online platforms that we investigated did not allow for quantifying search hits for different models or model-generated videos.
Subreddit selection: We selected 3 SFW and 3 NSFW subreddits dedicated to AI images and videos. For each category, we chose the most popular subreddits, all of which with over 100,000 members. The SFW subreddits that we selected were r/aiArt, r/StableDiffusion, r/aivideo. However, we omit the names of the NSFW subreddits to avoid driving internet traffic to them. However, readers can request the names with an email to us.
Model identification search terms: To identify instances of each video generation model across online platforms (Reddit, CivitAI, and the CivitAI archive), we used multiple search terms per model to account for different naming conventions, abbreviations, and variations used by the online community. These search terms were developed through preliminary exploration of how users reference these models in practice. For example, for Stable Video Diffusion, we searched for both the full name ”stable video diffusion” and the common abbreviation ”SVD.” Similarly, for Wan 2.x, we used both ”Wan” and ”Wan2” to capture different naming patterns. The complete list of search terms used for each model is shown in Table 2.
Table 2. Search terms used to identify models in online platforms. For each model, we used multiple search terms to capture different naming conventions and abbreviations used by the community.
Model Search Terms
Wan 2.x Wan, Wan2
HunyuanVideo Hunyuan, hunyuanvideo
Mochi 1 mochi, mochi-1
CogVideoX CogVideoX
LTX-Video LTXV, LTX-Video
Stable Video Diffusion SVD, stable video diffusion
Stable Virtual Camera SVC, stable virtual camera
AnimateDiff Lightning AnimateDiff, AnimateDiff-Lightning
Cosmos Cosmos
SeedVR2 SeedVR2
Stable Diffusion 1.x SD 1, SD 1.4, SD 1.5, stable diffusion 1
Stable Diffusion 2.x SD 2, SD 2.1, stable diffusion 2
Img2Vid and Framepacking tools: in Section 3.2, we restrict our analysis to AI video generators. However, it is important to note that a substantial fraction of AI videos online come from image generators scaffolded using tools such as Img2Vid and Framepack (zhang2025packing). However, these techniques are generally more time-consuming and typically produce lower-quality video than modern video generators.
The NSFW $\propto$ AIG-NCII assumption: In Section 3.2, we assume in our analysis that the uses of a model for NSFW content correlate with its uses for AIG-NCII. This is principally due to the ethical and legal hazards of searching for online AIG-NCII as well as the underground nature with which this content is generally made and distributed. We note that a flaw with this assumption relates to how some NSFW videos generated by AI systems are in a cartoon/anime style. For example, some models, such as PurpleSmart.AI’s Pony models, specialize in non-realistic forms of character art generation. However, from our exploration of videos generated with the models in Section 3.2, none appear to be disproportionately used for non-realistic NSFW character art compared to videorealistic NSFW content.